Please help me solve this problem!!
Find P(rolling 2 or 5) with one number cube.

Answer by rapaljer(4671)

There are a total of 6 equally likely possible outcomes.

P(2 or 5) = 2 chances out of 6 = 2/6 =1/3

Dr. Robert J. Rapalje, Retired
Professor of Mathematics
Seminole State College of Florida